OpenStudy (anonymous):

Which ordered pair is the point where the graph of y =1/2 x − 3 crosses the y-axis

OpenStudy (anonymous):

A. (−3, 0) B. (0, −3) C. (, 0) D. (3, )

OpenStudy (anonymous):

What is the equation of this relation? A. y = 2x + 2 B. y = 2x + 4 C. y = 2x − 2 D. y = 2x − 4

OpenStudy (anonymous):

perpendicular ?

OpenStudy (anonymous):

huh ?

OpenStudy (anonymous):

wait i got it wrong

OpenStudy (anonymous):

it wont let me enter the link

OpenStudy (anonymous):

the answers are What is the equation of this relation? A. y = 2x + 2 B. y = 2x + 4 C. y = 2x − 2 D. y = 2x − 4

OpenStudy (anonymous):

ok im a take a screenshot

OpenStudy (anonymous):

it is B

OpenStudy (anonymous):

A membership at a modern art museum costs $85 per year. Members pay a reduced price of $7.50 for entrance into each special exhibition. Which of the following statements is true of the annual cost of a membership at the museum? Let y = total cost and x = number of special exhibitions seen. A. The annual cost is shown by the linear function y = 85x + 7.50. The rate of change is 85, and the initial value is 7.50. B. The annual cost is shown by the linear function y = 92.50x. The rate of change is 92.50, and the initial value is 0. C. The annual cost is shown by the linear function y = 7.50x + 85. The rate of change is 7.50, and the initial value is 85. D. The annual cost is shown by the linear function y = x + 85. The rate of change is 1, and the initial value is 85

OpenStudy (anonymous):

2 more please

OpenStudy (anonymous):

C

OpenStudy (anonymous):

Using the graph below, which of the following equations best represents the amount of money Sara earns (m) in relation to the number of hours she works (h)? A. m = 15h B. m = 15 + h C. m = 30h D. m = 30 + h

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Which equation is graphed below? A. y = –2x – 2 B. y = 2x C. y = 2x + 2 D. y = 2x – 2

OpenStudy (anonymous):

D

OpenStudy (anonymous):

are u guessing ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

why do you think i am guessing? @t---_________---t

OpenStudy (anonymous):

just a little to quick

OpenStudy (anonymous):

all you havt to look at is where the segment hits the y axis and then look at the equations to see what number it is like if it hits at -2 on y axis then it has to end with -2 of the equation